# Low-probability Tokens Sustain Exploration in Reinforcement Learning with Verifiable Reward

### v1

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has propelled Large Language Models in complex reasoning, yet its scalability is often hindered by a training bottleneck where performance plateaus as policy entropy collapses, signaling a loss of exploration. Previous methods typically address this by maintaining high policy entropy, yet the precise mechanisms that govern meaningful exploration have remained underexplored. Our analysis suggests that an unselective focus on entropy risks amplifying irrelevant tokens and destabilizing training. This paper investigates the exploration dynamics within RLVR and identifies a key issue: the gradual elimination of valuable low-probability exploratory tokens, which we term reasoning sparks. We find that while abundant in pre-trained models, these sparks are systematically extinguished during RLVR due to over-penalization, leading to a degeneracy in exploration. To address this, we introduce Low-probability Regularization (Lp-Reg). Its core mechanism regularizes the policy towards a heuristic proxy distribution. This proxy is constructed by filtering out presumed noise tokens and re-normalizing the distribution over the remaining candidates. The result is a less-noisy proxy where the probability of reasoning sparks is amplified, which then serves as a soft regularization target to shield these valuable tokens from elimination via KL divergence. Experiments show that Lp-Reg enables stable on-policy training for around 1,000 steps, a regime where baseline entropy-control methods collapse. This sustained exploration leads to state-of-the-art performance, achieving a $60.17\%$ average accuracy on five math benchmarks, an improvement of $2.66\%$ over prior methods. Code is available at https://github.com/CarlanLark/Lp-Reg.
@@ 1 Introduction
The advent of large reasoning models has reshaped the trajectory of artificial intelligence, with paradigmatic examples including OpenAI O1 (OpenAI et al., 2024) and DeepSeek-R1 (DeepSeek-AI et al., 2025). A central technique underpinning these systems is reinforcement learning with verifiable reward (RLVR), which assigns reward to verifiable solutions through rule-based verification. These models generate extended chain-of-thought (CoT) reasoning (Wei et al., 2023) to solve challenging problems in domains like mathematical olympiads (He et al., 2024b). However, a notable bottleneck emerges during RL training that limits its scalability, frequently culminating in a performance plateau and subsequent collapse. This failure is consistently accompanied by a rapid decay in policy entropy, indicating a severe loss of exploration capacity (Yu et al., 2025; Cui et al., 2025; Wang et al., 2025).
Previous approaches have recognized this declining exploration, attempting to address it through various entropy control mechanisms. Methods such as adaptive entropy regularization (He et al., 2025), high entropy change blocking (Cui et al., 2025), or selective token updates (Wang et al., 2025) aim to maintain higher entropy as a proxy for exploration. However, relying on overall entropy can be an indirect and imprecise tool. An indiscriminate focus on maximizing randomness risks amplifying noise and destabilizing training (Ömer Veysel Çağatan & Akgün, 2025), suggesting a deeper issue beyond simply the quantity of randomness.
@@
Figure 1: Selectively preserving low-probability tokens is key to overcoming performance plateaus in reasoning RL. (a) An illustration of a reasoning spark. (b) Standard GRPO training reaches a performance plateau and collapses, accompanied by decaying entropy. An indiscriminate entropy bonus (GRPO + Entropy Loss) leads to an even faster collapse. (c) We reveal the cause: GRPO systematically suppresses the low-probability sampling of important exploratory tokens (like “wait”), and forces these tokens’ sampling distributions to collapse towards high probabilities. Ent-Loss fails to fix this. In contrast, our method, Lp-Reg, successfully preserves a healthy, wide distribution, sustaining exploration. (d) The failure of entropy bonuses is explained by their indiscriminate nature: they amplify the low-probability sampling of irrelevant tokens, creating noise and thereby degrading exploration quality. The aggregated statistics in (c) and (d) demonstrate a systemic effect beyond single-token instances. Detailed plots for individual tokens are available in Appendix B.1.
@@
Our experimental evaluation demonstrates the effectiveness of Lp-Reg. Our method enables stable on-policy training for around 1,000 steps, a regime where many entropy-control methods have collapsed, resulting in better performance. On five widely used math benchmarks, this results in a $60.17\%$ average accuracy on Qwen3-14B-Base, improving upon prior methods by $2.66\%$. Our contributions are summarized as follows:
@@
  • We demonstrate through extensive experiments that Lp-Reg achieves state-of-the-art performance, while also enabling stable on-policy training over extended periods where baselines collapse.
@@ 2 Related Work
Recently, reinforcement learning has become the dominant framework for enhancing the reasoning abilities of large language models (LLMs) (OpenAI et al., 2024; DeepSeek-AI et al., 2025). By leveraging automatic checkers or symbolic verification, reinforcement learning with verifiable rewards (RLVR) achieved further breakthroughs in improving the reasoning capability of LLMs (Shao et al., 2024a; Yang et al., 2025; Team et al., 2025). Based on RLVR and GRPO (Shao et al., 2024a), subsequent methods such as DAPO Yu et al. (2025), VAPO (Yue et al., 2025), and other policy optimization variants (Zhao et al., 2025; Cui et al., 2025; Zheng et al., 2025) have been proposed to improve the stability, efficiency, and scalability of RL for reasoning models.
@@
A recurring difficulty in training reasoning models with RL is the rapid collapse of policy entropy during the early stages of training. This phenomenon, which reflects excessive exploitation and insufficient exploration, has been widely recognized as a bottleneck for scaling RL in reasoning models. To mitigate collapse, researchers have explored several directions, including selectively regularizing updates at high-entropy “forking” tokens (Wang et al., 2025), amplifying advantages at exploratory positions (Cheng et al., 2025), modifying clipping strategies (Yu et al., 2025; Zhao et al., 2025; Cui et al., 2025; Zheng et al., 2025), or doing weight clipping (MiniMax et al., 2025; Su et al., 2025). While these methods primarily operate by monitoring policy entropy, which is correlational rather than causal to exploration, our analysis delves directly into the next-token prediction distribution. This allows for a more semantically grounded and causally-informed investigation of the probabilities of individual candidates and their role in exploration dynamics.
@@
As the capabilities of Large Language Models (LLMs) have rapidly advanced, they have demonstrated an increasingly strong and reliable sense of intrinsic confidence (Saurav et al., 2022; Loka et al., 2024; Amir et al., 2025). Research investigates how these intrinsic confidence signals, often reflected in the next-token prediction distribution, can guide complex reasoning and exploration (Amirhosein et al., 2025; Xuezhi & Denny, 2024; Xuandong et al., 2025). Studies have shown that tokens with higher relative probabilities in the next-token prediction are often more contextually appropriate than their lower-probability counterparts (Nguyen et al., 2025; Xu et al., 2025; Fu et al., 2025). Building on this, some work has explored entropy minimization, which sharpens the model’s confidence distribution. This approach can improve inference performance by encouraging the model to commit to consistent and high-confidence solution paths (Gao et al., 2025; Agarwal et al., 2025). Our work builds upon a similar insight, leveraging the model’s intrinsic confidence to distinguish between valuable reasoning sparks and irrelevant noise within this low-probability range.
@@ 3.1 Reinforcement Learning with Verifiable Rewards
Reinforcement learning (RL) has played a critical role in LLMs (Murphy, 2024). Formally,
@@
where $r(o,a)$ denotes the reward assigned to an output $o$ given a reference answer $a$. In reinforcement learning with verifiable rewards (RLVR), this reward is computed through rule-based functions, such as Math-Verify [^1] https://github.com/huggingface/Math-Verify. Recent studies have demonstrated that large-scale RLVR encourages models to perform more deliberative reasoning by producing extended chains of thought prior to the final prediction, thereby substantially improving their capacity to solve complex problems (DeepSeek-AI et al., 2025). In practice, Eq. 1 is typically optimized using policy gradient methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024b).
@@ 4.1 Proxy Distribution $\pi_{\text{proxy}}$
    • Fixed threshold: A simple approach where $\tau$ is a constant hyperparameter, e.g., $0.02$.
    • Min-p threshold: Following (Nguyen et al., 2025), $\tau$ is defined relative to the peak probability: $\tau=\kappa\cdot\max_{o^{\prime}\in V}\pi_{\bm{\theta}}(o^{\prime}|\cdot)$, where $\kappa\in(0,1)$ is a hyperparameter. This makes the filtering adaptive to the distribution’s sharpness.
@@ 4.2 Low-probability Regularization Objective
$J(\bm{\theta})=$ $\mathbb{E}_{\mathcal{B}\sim\mathcal{D},(q,a)\sim\mathcal{B},\{o_{i}\}_{i=1}^{G}\sim\pi_{\bm{\theta}_{\text{old}}}(\cdot|q)}\Bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Big[\mathrm{clip}(r_{i,t}(\bm{\theta}),0,U)\cdot A_{i,t}$ (6)
$-\beta\cdot\mathbb{I}\left[\pi_{\bm{\theta}}(o_{i,t}|q,o_{i,<t})<\delta_{\rho}^{\mathcal{B}}\land\pi_{\text{proxy}}(o_{i,t}|q,o_{i,<t})>0\land A_{i,t}<0\right]$
$\qquad\cdot\mathcal{D}_{\mathrm{KL}}\left(\pi_{\text{proxy}}(\cdot|q,o_{i,<t})\|\pi_{\bm{\theta}}(\cdot|q,o_{i,<t})\right)\Big]\Bigg]$
@@
The second term is the Lp-Reg penalty, designed to protect reasoning sparks. It is activated by the indicator function $\mathbb{I}[\cdot]$ only for tokens that satisfy three conditions simultaneously: first, their sampling probability $\pi_{\bm{\theta}}$ is below a dynamic low-percentile threshold $\delta_{\rho}^{\mathcal{B}}$, which is calculated as the lowest $\rho$-th percentile of the sampling probabilities of all tokens within the current training batch $\mathcal{B}$; second, their probability in the proxy distribution $\pi_{\text{proxy}}$ is greater than zero; and third, the token receives a negative advantage signal ($A_{i,t}<0$). This final condition ensures the regularization applies exclusively to tokens receiving a negative learning signal, preventing their potential over-penalization while leaving updates from positive experiences unaffected.
@@ 5.1 Experimental Setup
We compare Lp-Reg against a suite of strong baselines, including a foundational algorithm and several state-of-the-art methods designed to enhance exploration through entropy control. Our primary baseline is GRPO (Shao et al., 2024a), a value-free policy optimization algorithm that employs group-relative advantage estimation, making it a common choice for RLVR. To represent classical entropy regularization methods, we implement GRPO + Entropy Loss, which directly incorporates the principles of Maximum Entropy RL by adding a policy entropy bonus to the GRPO objective function. We also compare against several advanced methods: Clip-Higher (Yu et al., 2025), a core component of DAPO that encourages higher entropy by using an asymmetric clipping range in the PPO objective; Selective High-Entropy Training (80/20) (Wang et al., 2025), a method that restricts policy gradient updates to only the top 20% of tokens with the highest generation entropy; KL-Cov (Cui et al., 2025), which prevents entropy collapse by applying a selective KL-divergence penalty to tokens with the highest covariance between their log probabilities and advantages; and GSPO (Zheng et al., 2025), which modifies the clipping mechanism to operate at the sequence level to promote higher training entropy.
Training Settings
@@
All experiments are conducted within the verl (Sheng et al., 2024) framework to ensure a standardized and fair comparison. We use two base models for our experiments: Qwen3-14B-Base and Qwen2.5-32B. The 14B models are trained on 32 NVIDIA H20 GPUs, while the 32B model is trained on 64 NVIDIA H20 GPUs, with an average training step time of 6001400600-1400 seconds. For the reinforcement learning from verifier rewards (RLVR) phase, models are trained on the Dapo-Math-17K Yu et al. (2025) dataset with a maximum response length of $8,192$ tokens. We use a global batch size of 256. For off-policy methods, we use a mini-batch size of 32, resulting in 8 gradient updates per rollout. To ensure a fair comparison, a “step” in our experimental results consistently refers to a single rollout for all methods. Consequently, each reported step for off-policy training corresponds to 8 gradient updates. A constant learning rate of $1\times 10^{-6}$ is applied without a warmup schedule. We set the group number as $8$ for all GRPO-based methods. To ensure numerical stability, we set the policy gradient’s clipping by setting the upper bound of the importance sampling ratio to $U=10$. For our proposed Lp-Reg, which uses the min-p threshold, we set the probability percentile threshold $\rho$ to $0.5\%$ for Qwen2.5-32B-Base and $1\%$ for Qwen3-14B-Base, the KL regularization coefficient $\beta$ to $1.0$, and the min-p ratio $\kappa$ to $0.02$. The proxy distribution, $\pi_{\text{proxy}}$, is constructed from the data-generating policy ($\pi_{\theta_{\text{old}}}$ in the off-policy setting and the current policy $\pi_{\bm{\theta}}$ in the on-policy setting). For all baseline methods, we adopt the hyperparameters specified in their original public implementations to ensure a faithful reproduction. Specifically for the GRPO + Entropy Loss baseline, we set the entropy coefficient to $0.002$ within the verl framework.
Evaluation

### v2

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has propelled Large Language Models in complex reasoning, yet its scalability is often hindered by a training bottleneck where performance plateaus as policy entropy collapses, signaling a loss of exploration. Previous methods typically address this by maintaining high policy entropy, yet the precise mechanisms that govern meaningful exploration have remained underexplored. Our analysis suggests that an unselective focus on entropy risks amplifying irrelevant tokens and destabilizing training. This paper investigates the exploration dynamics within RLVR and identifies a key issue: the gradual elimination of valuable low-probability exploratory tokens, which we term reasoning sparks. We find that while abundant in pre-trained models, these sparks are systematically extinguished during RLVR due to over-penalization, leading to a degeneracy in exploration. To address this, we introduce Low-probability Regularization (Lp-Reg). Its core mechanism regularizes the policy towards a heuristic proxy distribution. This proxy is constructed by filtering out presumed noise tokens and re-normalizing the distribution over the remaining candidates. The result is a less-noisy proxy where the probability of reasoning sparks is amplified, which then serves as a soft regularization target to shield these valuable tokens from elimination via KL divergence. Experiments show that Lp-Reg enables stable on-policy RL, sustaining continuous scaling across $3,000$ training steps and $81,204$ GPU-hours, where baseline entropy-control methods collapse. This sustained exploration leads to state-of-the-art performance, achieving a $60.17\%$ average accuracy on five math benchmarks, an improvement of $2.66\%$ over prior methods. Code is available at https://github.com/CarlanLark/Lp-Reg.
@@ 1 Introduction
The advent of large reasoning models has reshaped the trajectory of artificial intelligence, with paradigmatic examples including OpenAI O1 (openai2024openaio1card) and DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability). A central technique underpinning these systems is reinforcement learning with verifiable reward (RLVR), which assigns reward to verifiable solutions through rule-based verification. These models generate extended chain-of-thought (CoT) reasoning (wei2023chainofthoughtpromptingelicitsreasoning) to solve challenging problems in domains like mathematical olympiads (he2024olympiadbench). However, a notable bottleneck emerges during RL training that limits its scalability, frequently culminating in a performance plateau and subsequent collapse. This failure is consistently accompanied by a rapid decay in policy entropy, indicating a severe loss of exploration capacity (yu2025dapoopensourcellmreinforcement; cui2025entropymechanismreinforcementlearning; wang20258020rulehighentropyminority).
Previous approaches have recognized this declining exploration, attempting to address it through various entropy control mechanisms. Methods such as adaptive entropy regularization (he2025skyworkopenreasoner1), high entropy change blocking (cui2025entropymechanismreinforcementlearning), or selective token updates (wang20258020rulehighentropyminority) aim to maintain higher entropy as a proxy for exploration. However, relying on overall entropy can be an indirect and imprecise tool. An indiscriminate focus on maximizing randomness risks amplifying noise and destabilizing training (2025failuremodesmaximumentropy), suggesting a deeper issue beyond simply the quantity of randomness.
@@
Figure 1: Selectively preserving low-probability tokens is key to overcoming performance plateaus in reasoning RL. (a) An illustration of a reasoning spark. (b) Standard GRPO training reaches a performance plateau and collapses, accompanied by decaying entropy. An indiscriminate entropy bonus (GRPO + Entropy Loss) leads to an even faster collapse. (c) We reveal the cause: GRPO systematically suppresses the low-probability sampling of important exploratory tokens (like “wait”), and forces these tokens’ sampling distributions to collapse towards high probabilities. Entropy Loss fails to fix this. In contrast, our method, Lp-Reg, successfully preserves a healthy, wide distribution, sustaining exploration. (d) The failure of entropy bonuses is explained by amplifying the low-probability sampling of irrelevant tokens, creating noise, and thereby degrading exploration quality. The aggregated statistics in (c) and (d) demonstrate a systemic effect beyond single-token instances. Detailed plots for individual tokens are available in Appendix B.1.
@@
Our experimental evaluation demonstrates the effectiveness of Lp-Reg. Our method enables stable on-policy training for around $3,000$ steps over $81,204$ GPU-hours, a regime where many entropy-control methods have collapsed, resulting in better performance. On five widely used math benchmarks, this results in a $60.17\%$ average accuracy on Qwen3-14B-Base, improving upon prior methods by $2.66\%$. Our contributions are summarized as follows:
@@
  • We demonstrate through extensive experiments that Lp-Reg achieves state-of-the-art performance, while also enabling stable on-policy continuous scaling over extended periods where baselines collapse.
@@ 2 Related Work
Recently, reinforcement learning has become the dominant framework for enhancing the reasoning abilities of large language models (LLMs) (openai2024openaio1card; deepseekai2025deepseekr1incentivizingreasoningcapability). By leveraging automatic checkers or symbolic verification, reinforcement learning with verifiable rewards (RLVR) achieved further breakthroughs in improving the reasoning capability of LLMs (shao2024deepseekmathpushinglimitsmathematical; yang2025qwen3technicalreport; kimiteam2025kimik15scalingreinforcement). Based on RLVR and GRPO (shao2024deepseekmathpushinglimitsmathematical), subsequent methods such as DAPO yu2025dapoopensourcellmreinforcement, VAPO (yue2025vapoefficientreliablereinforcement), and other policy optimization variants (zhao2025geometricmeanpolicyoptimization; cui2025entropymechanismreinforcementlearning; zheng2025groupsequencepolicyoptimization) have been proposed to improve the stability, efficiency, and scalability of RL for reasoning models.
@@
A recurring difficulty in training reasoning models with RL is the rapid collapse of policy entropy during the early stages of training. This phenomenon, which reflects excessive exploitation and insufficient exploration, has been widely recognized as a bottleneck for scaling RL in reasoning models. To mitigate collapse, researchers have explored several directions, including selectively regularizing updates at high-entropy “forking” tokens (wang20258020rulehighentropyminority), amplifying advantages at exploratory positions (cheng2025reasoningexplorationentropyperspective), modifying clipping strategies (yu2025dapoopensourcellmreinforcement; zhao2025geometricmeanpolicyoptimization; cui2025entropymechanismreinforcementlearning; zheng2025groupsequencepolicyoptimization), or doing weight clipping (minimax2025minimaxm1scalingtesttimecompute; su2025klearreasoneradvancingreasoningcapability). While these methods primarily operate by monitoring policy entropy, which is correlational rather than causal to exploration, our analysis delves directly into the next-token prediction distribution. This allows for a more semantically grounded and causally-informed investigation of the probabilities of individual candidates and their role in exploration dynamics.
@@
As the capabilities of Large Language Models (LLMs) have rapidly advanced, they have demonstrated an increasingly strong and reliable sense of intrinsic confidence (saurav2022language; loka2024confidence; amir2025confidence). Research investigates how these intrinsic confidence signals, often reflected in the next-token prediction distribution, can guide complex reasoning and exploration (amirhosein2025guided; xuezhi2024chainofthought; xuandong2025learning). Studies have shown that tokens with higher relative probabilities in the next-token prediction are often more contextually appropriate than their lower-probability counterparts (nguyen2025turningheatminpsampling; xu2025adaptiveterminationmultiroundparallel; fu2025deepthinkconfidence). Building on this, some work has explored entropy minimization, which sharpens the model’s confidence distribution. This approach can improve inference performance by encouraging the model to commit to consistent and high-confidence solution paths (gao2025oneshotentropyminimization; agarwal2025unreasonableeffectivenessentropyminimization). Our work builds upon a similar insight, leveraging the model’s intrinsic confidence to distinguish between valuable reasoning sparks and irrelevant noise within this low-probability range.
@@ 3.1 Reinforcement Learning with Verifiable Rewards
Reinforcement learning (RL) has played a critical role in LLMs (murphy2024reinforcement). Formally,
@@
where $r(o,a)$ denotes the reward assigned to an output $o$ given a reference answer $a$. In reinforcement learning with verifiable rewards (RLVR), this reward is computed through rule-based functions, such as Math-Verify [^1] https://github.com/huggingface/Math-Verify. Recent studies have demonstrated that large-scale RLVR encourages models to perform more deliberative reasoning by producing extended chains of thought prior to the final prediction, thereby substantially improving their capacity to solve complex problems (deepseekai2025deepseekr1incentivizingreasoningcapability). In practice, Eq. 1 is typically optimized using policy gradient methods, such as Proximal Policy Optimization (PPO) (schulman2017proximal) and Group Relative Policy Optimization (GRPO) (shao2024deepseekmath).
@@ 4.1 Proxy Distribution $\pi_{\text{proxy}}$
    • Fixed threshold: A simple approach where $\tau$ is a constant hyperparameter, e.g., $\tau=0.02$.
    • Min-p threshold: Following (nguyen2025turningheatminpsampling), $\tau$ is defined relative to the peak probability: $\tau=\kappa\cdot\max_{o^{\prime}\in V}\pi_{\bm{\theta}}(o^{\prime}|\cdot)$, where $\kappa\in(0,1)$ is a hyperparameter. This makes the filtering adaptive to the distribution’s sharpness.
@@ 4.2 Low-probability Regularization Objective
$J_{\text{Lp-Reg}}(\bm{\theta})=$ $\mathbb{E}_{\mathcal{B}\sim\mathcal{D},(q,a)\sim\mathcal{B},\{o_{i}\}_{i=1}^{G}\sim\pi_{\bm{\theta}_{\text{old}}}(\cdot|q)}\Bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\bigg[\mathrm{clip}(r_{i,t}(\bm{\theta}),0,U)\cdot A_{i,t}$ (6)
$-\beta\cdot\mathbb{I}\bigg[\underbrace{\pi_{\bm{\theta}}(o_{i,t}|q,o_{i,<t})<\delta_{\rho}^{\mathcal{B}}}_{\text{low probability}}\land\underbrace{\pi_{\text{proxy}}(o_{i,t}|q,o_{i,<t})>0}_{\text{not noise}}\land\underbrace{A_{i,t}<0}_{\text{neg. sample}}\bigg]$
$\qquad\cdot\mathcal{D}_{\mathrm{KL}}\left(\pi_{\text{proxy}}(\cdot|q,o_{i,<t})\|\pi_{\bm{\theta}}(\cdot|q,o_{i,<t})\right)\bigg]\Bigg]$
@@
The second term is the Lp-Reg penalty, designed to protect reasoning sparks. It is activated by the indicator function $\mathbb{I}[\cdot]$ only for tokens that satisfy three conditions simultaneously: first, their sampling probability $\pi_{\bm{\theta}}$ is below a dynamic low-percentile threshold $\delta_{\rho}^{\mathcal{B}}$, which is calculated as the lowest $\rho$-th percentile of the sampling probabilities of all tokens within the current training batch $\mathcal{B}$; second, their probability in the proxy distribution $\pi_{\text{proxy}}$ is greater than zero, means $o_{i,t}$ is not a noise token; and third, the token receives a negative advantage signal ($A_{i,t}<0$). This final condition ensures the regularization applies exclusively to tokens receiving a negative learning signal, preventing their potential over-penalization while leaving updates from positive experiences unaffected. We ablate all three conditions in Section 5.3.
@@ 5.1.1 Baselines
We compare Lp-Reg against a suite of strong baselines, including a foundational algorithm and several state-of-the-art methods designed to enhance exploration through entropy control. Our primary baseline is GRPO (shao2024deepseekmathpushinglimitsmathematical), a value-free policy optimization algorithm that employs group-relative advantage estimation, making it a common choice for RLVR. To represent classical entropy regularization methods, we implement GRPO + Entropy Loss, which directly incorporates the principles of Maximum Entropy RL by adding a policy entropy bonus to the GRPO objective function. We also compare against several advanced methods: Clip-Higher (yu2025dapoopensourcellmreinforcement), a core component of DAPO that encourages higher entropy by using an asymmetric clipping range in the PPO objective; Selective High-Entropy Training (80/20) (wang20258020rulehighentropyminority), a method that restricts policy gradient updates to only the top 20% of tokens with the highest generation entropy; KL-Cov (cui2025entropymechanismreinforcementlearning), which prevents entropy collapse by applying a selective KL-divergence penalty to tokens with the highest covariance between their log probabilities and advantages; and GSPO (zheng2025groupsequencepolicyoptimization), which modifies the clipping mechanism to operate at the sequence level to promote higher training entropy.
5.1.2 Training Settings
All experiments are conducted within the verl (sheng2024hybridflow) framework to ensure a standardized and fair comparison. To mitigate unfair comparisons arising from different convergence speeds during the early stages of training, we trained models until their performance saturated. This approach ensures a more stable and equitable evaluation point. Consequently, our comparisons are based on approximately $1,000$ training steps for the Qwen3-14B-Base model and $800$ for the Qwen2.5-32B model. Each training requires about $8,000$ GPU hours on 32 NVIDIA H20 GPUs for the 14B model and $16,000$ GPU hours on 64 NVIDIA H20 GPUs for the 32B model. However, for models that experienced a training collapse, which we define as a performance drop greater than $10\%$ in accuracy, we implemented early stopping to conserve computational resources. To assess whether low-probability tokens sustain exploration in RLVR, we further trained the Qwen2.5-32B model for $3,000$ steps over $81,204$ GPU-hours with our Lp-Reg and evaluated its training stability.
@@ 5.1.2 Training Settings
For the reinforcement learning from verifier rewards (RLVR) phase, models are trained on the Dapo-Math-17K (yu2025dapoopensourcellmreinforcement) dataset with a maximum response length of $8,192$ tokens. We use a global batch size of 256. For off-policy methods, we use a mini-batch size of 32, resulting in 8 gradient updates per rollout. To ensure a fair comparison, a “step” in our experimental results consistently refers to a single rollout for all methods. Consequently, each reported step for off-policy training corresponds to 8 gradient updates. A constant learning rate of $1\times 10^{-6}$ is applied without a warmup schedule. We set the group number as $8$ for all GRPO-based methods. To ensure numerical stability, we set the policy gradient’s clipping by setting the upper bound of the importance sampling ratio to $U=10$. For our proposed Lp-Reg, which uses the min-p threshold, we set the probability percentile threshold $\rho$ to $0.5\%$ for Qwen2.5-32B-Base and $1\%$ for Qwen3-14B-Base, the KL regularization coefficient $\beta$ to $1.0$, and the min-p ratio $\kappa$ to $0.02$. The proxy distribution, $\pi_{\text{proxy}}$, is constructed from the data-generating policy ($\pi_{\theta_{\text{old}}}$ in the off-policy setting and the current policy $\pi_{\bm{\theta}}$ in the on-policy setting). For all baseline methods, we adopt the hyperparameters specified in their original public implementations to ensure a faithful reproduction. Specifically for the GRPO + Entropy Loss baseline, we set the entropy coefficient to $0.002$ within the verl framework.
5.1.3 Evaluation